# Context-Driven Performance Modeling for Causal Inference Operators on Neural Processing Units

### v1

@@ III-B Device Utilization Analysis
To understand the performance bottlenecks, we profile the execution time spent on the NPU’s primary components: the DPU, DMA, and SHAVE cores. Table II presents the utilization breakdown for Fourier State-Space Attention (FSA) and Decayed Recurrent Attention (DRA), which serves as a proxy for retentive decay mechanisms.
@@
TABLE II: Device Utilization Breakdown (%). At long contexts, FSA becomes DMA-bound while DRA becomes SHAVE-bound.
Model Context DPU (%) DMA (%) SHAVE (%) Bottleneck
Fourier 128 56.4 23.1 20.5 DPU
256 60.8 25.3 13.9 DPU
512 47.2 46.9 5.9 DMA / DPU
1024 46.6 48.9 4.5 DMA
2048 46.2 52.5 1.2 DMA
4096 48.4 51.3 0.3 DMA
8192 61.1 38.9 0.0 DPU
Retentive 128 68.4 0.0 31.6 DPU
256 64.9 0.0 35.1 DPU
512 61.9 0.0 38.1 DPU
1024 34.9 0.0 65.1 SHAVE
2048 24.6 0.0 75.4 SHAVE
4096 28.1 0.0 71.9 SHAVE
8192 23.6 0.0 76.4 SHAVE
Figure 4 illustrates how NPU utilization shifts with context length. Fourier transitions from DPU- to DMA-bound, while Retentive becomes increasingly SHAVE-bound due to element-wise $\operatorname{softmax}$ overhead.
Figure 4: Utilization of NPU subcomponents (DPU, DMA, SHAVE) for Fourier and Retentive attention mechanisms as a function of context length (log scale). Fourier transitions from DPU-bound to DMA-bound at longer contexts, while Retentive attention becomes increasingly SHAVE-bound due to $\operatorname{softmax}$ and element-wise operations.
III-C Scaling of Latency with Context Length
We measure the end-to-end latency of four attention mechanisms—Fourier Structured Attention (FSA), DRA, Toeplitz Structured Attention (TSA), and Causal Linear Attention (CLA)—across a range of context lengths from 128 to 8192. The results, summarized in Table III, reveal distinct scaling properties for each model.
TABLE III: Latency scaling (ms) as a function of context length for four attention operators.
Context Length Fourier Retentive Toeplitz Linear
128 0.39 0.19 0.06 0.09
256 0.79 0.37 0.08 0.13
512 2.54 0.97 0.11 0.24
1024 6.50 2.52 0.18 0.44
2048 15.24 10.04 0.35 0.72
4096 45.69 39.52 0.59 1.52
8192 347.79 85.41 1.01 3.16
@@ III-C Scaling of Latency with Context Length
Figure 5: Latency scaling of four causal inference operators with respect to context length (log scale). Toeplitz and Linear exhibit near-linear latency growth, while Fourier scales poorly at long contexts due to DPU and DMA bottlenecks.
III-D Performance Scaling and Bottleneck Analysis
We analyze the performance of each kernel by examining latency, throughput, pipeline efficiency, and memory access patterns across varying context lengths. The results, summarized in Tables IV and V, reveal significant architectural bottlenecks for quadratic-time algorithms when deployed on the NPU.
As shown in Table IV, the latency of standard Causal Attention Masking scales quadratically with the sequence length ($N$). This leads to a dramatic drop in throughput, with Causal processing only 4 operations per second at a context of 8192. In contrast, sub-quadratic methods like TSA and Linear Attention maintain low latency and high throughput, demonstrating their suitability for long-context applications on this hardware.
TABLE IV: Latency and throughput scaling at short ($N=512$) and long ($N=8192$) contexts.
@@ III-D Performance Scaling and Bottleneck Analysis
The underlying cause of this performance degradation is revealed in Table V. At a context length of 8192, Causal and Retentive attention mechanisms suffer from extremely high pipeline stall rates (96.7% and 94.8%, respectively). This indicates that the NPU pipeline is mostly idle, waiting for data to be fetched from memory. The pull stage, responsible for data retrieval, is the primary contributor to these stalls. This is further corroborated by the poor cache efficiency scores for these models (7.7% for Causal and 28.1% for Retentive), which signify a low degree of data reuse and frequent, costly access to main memory. In contrast, the more structured access patterns of Linear and Toeplitz attention lead to much higher cache efficiency and significantly fewer pipeline stalls.
TABLE V: Efficiency metrics at long context lengths. Stall and cache values are percentages; reuse is in milliseconds.
@@
III-E Impact of State Dimension
To assess the impact of model parameters on performance, we benchmarked several kernels at a fixed context length ($N=4096$) while increasing the state dimension ($d_{\text{state}}$) from the default of 16 to 128. As detailed in Table VI, increasing the state dimension leads to a predictable rise in latency across all tested models. Notably, Fourier is the most sensitive, with its latency increasing by over $10\times$, highlighting its high computational cost with respect to the state size. Linear and Toeplitz also show increased latency but remain significantly more efficient, confirming that their performance is a function of both sequence length and model dimension.
TABLE VI: Latency impact of increasing state dimension ($d_{\text{state}}$) at fixed context length $N=4096$.
@@ IV-B Operator Intensity Characterization
TABLE VII: Operational intensity and measured performance at $N=4096$, $d_{h}=64$ (16-bit precision).
@@ IV-C Roofline Analysis
$10^{1}$$10^{1.3}$$10^{1.7}$$10^{2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$10^{2.7}$$10^{2.19}$Causal (4.3%)Linear (27.3%)Fourier (0.7%)Retentive (33.4%)Toeplitz (15.2%)Operational Intensity II [Ops/Byte]Performance $P$ [GOP/s]Roofline Model with Effective CeilingsEffective Memory CeilingEffective Compute Ceiling
Figure 7: Roofline model revealing severe hardware under utilization. Quadratic operators achieve <<5% of their compute ceiling despite high intensity, while linear methods approach 27% of their memory-bound limit.
@@ IV-E Efficiency Analysis
TABLE VIII: Hardware utilization metrics at $N=4096$.
@@
This model proves that memory access patterns, not theoretical FLOP counts, dominate NPU performance. Operators must be co-designed with: (1) systolic-compatible dataflow, (2) predictable memory access, and (3) minimized DMA transfers to approach effective hardware limits.
@@ V Discussion
SHAVE Core Bottlenecks in Element-wise Operations While NPUs excel at matrix multiplication via systolic arrays, element-wise operations (e.g., $\operatorname{softmax}$, $\operatorname{scaling}$) execute on general-purpose SHAVE cores. At $N>1024$, these operations dominate latency in recurrent attention variants (up to 76% utilization). Kernel fusion or dedicated acceleration is essential.
DMA Management for Memory-Intensive Ops Tensor concatenation and state management consume 40-50% of cycles in Fourier attention (Table II). DMA overheads stem from frequent allocation/deallocation of large buffers. Offloading these operations to the CPU reduces latency by 32% in our tests.
@@ VI-A Long-Context Inference on Edge Platforms
Prior work has explored deploying transformer-based causal large language models (LLMs) on edge platforms [13], including hardware-specific optimizations for ARM CPUs [14] and FPGA-based execution through frameworks like llama.cpp [15, 16]. While these efforts target on-device inference, they are not designed for long-context scenarios and do not address the associated memory and compute bottlenecks that emerge in attention-heavy models. Our work addresses this gap by empirically analyzing a range of causal inference mechanisms—including standard transformers and structured state-space models (SSMs)—under long-context settings. This enables us to derive architectural insights that inform the co-design of attention mechanisms for Neural Processing Units (NPUs), supporting more efficient hardware-aware deployment strategies.
@@ VI-B Acceleration of Sequence Models on NPUs
Several efforts have investigated transformer acceleration on NPUs [17, 18], typically through operator-level scheduling or compiler-level block partitioning. However, these approaches fall short in capturing the fine-grained resource behavior required for efficient long-context inference. Other work has focused on optimizing SSMs for NPUs [19, 20], leveraging architectural properties such as linear recurrence and memory compression. While successful within their respective domains, these strategies are not directly applicable to transformer-style causal attention. In contrast, our approach uses execution profiling and performance modeling—grounded in structured operator variants—to analyze architectural trade-offs across attention and SSM-style models. By leveraging structured state-space duality (SSD), we characterize how causal operators interact with NPU memory and compute hierarchies, enabling more informed co-design for future inference systems.
@@ VII Conclusion
Deploying long-context AI on edge devices centers on resolving the fundamental mismatch between NPU architectures—optimized for dense, regular computations—and the memory-intensive, irregular access patterns of quadratic attention. Our analysis reveals catastrophic hardware underutilization (>>95% stalls) in standard approaches, while demonstrating that structured sub-quadratic operators (Toeplitz, Linear) transform the bottleneck into manageable bandwidth constraints. This necessitates a paradigm shift: throughput gains come not from incremental attention optimizations, but from co-designing causal operators that respect systolic dataflows and memory hierarchies. By aligning algorithmic structure with NPU execution models, we unlock the path to pervasive, private, and powerful edge AI.

### v2

@@ III-E Device Utilization Analysis
To understand the performance bottlenecks, we profile the execution time spent on the NPU’s primary components: the DPU, DMA, and SHAVE cores. Table III presents the utilization breakdown for Fourier State-Space Attention (FSA) and Decayed Recurrent Attention (DRA), which serves as a proxy for retentive decay mechanisms.
@@
TABLE III: Device utilization breakdown (%) across operators and context lengths. Conv1D Direct/Dilated achieve near-perfect DPU utilization, while Conv1D FFT exhibits catastrophic SHAVE-boundedness.
Operator Context DPU (%) DMA (%) SHAVE (%) Bottleneck
Attention Variants
Full Causal 8192 4.3 27 68.7 SHAVE
Retentive 8192 23.6 0.0 76.4 SHAVE
Fourier 8192 61.1 38.9 0.0 DPU
Toeplitz 4096 70.0 30.0 0.0 DPU
Semiseparable 4096 85.9 0.9 13.2 DPU
State Space Models
SSM Sequential 1024 100.0 0.0 0.0 DPU
SSM Parallel 4096 54.8 9.2 35.9 DPU
SSM Chunked 1024 94.7 5.3 0.0 DPU
1D Convolutions
Conv1D Direct 4096 99.6 0.4 0.0 DPU
Conv1D Dilated 4096 99.6 0.4 0.0 DPU
Conv1D FFT 512 41.0 0.2 58.8 SHAVE
Table III illustrates how NPU utilization shifts with context length. Fourier transitions from DPU- to DMA-bound, while Retentive becomes increasingly SHAVE-bound due to element-wise $\operatorname{softmax}$ overhead.
III-F Scaling of Latency with Context Length
We measure the end-to-end latency of four attention mechanisms—Fourier Structured Attention (FSA), DRA, Toeplitz Structured Attention (TSA), and Causal Linear Attention (CLA)—across a range of context lengths from 128 to 8192. The results, summarized in Table IV, reveal distinct scaling properties for each model.
TABLE IV: Latency scaling (ms) as a function of context length for eleven causal operators on NPU. Conv1D Direct/Dilated and SSM variants demonstrate superior efficiency, while Conv1D FFT exhibits catastrophic SHAVE-bound performance.
Operator N=128 N=512 N=1024 N=2048 N=4096 N=8192
Attention Variants
Full Causal 2.43 5.47 18.23 29.26 70.43 305.23
Fourier 32.19 111.47 - - - -
Retentive 0.78 2.46 4.70 10.05 26.66 72.53
Toeplitz 1.59 1.88 2.64 5.02 7.54 13.36
Semiseparable 0.78 2.22 4.28 10.59 26.87 79.33
State Space Models
SSM Sequential 4.24 16.30 26.38
SSM Parallel 0.69 2.71 2.53 3.88 7.75 13.11
SSM Chunked 6.16 22.31 30.35 57.91
1D Convolutions
Conv1D Direct 2.03 8.74 24.64 31.85 54.71
Conv1D Dilated 3.07 9.68 23.69 37.48 74.85
Conv1D FFT 28.00 304.64
@@ III-F Scaling of Latency with Context Length
III-G Performance Scaling and Bottleneck Analysis
We analyze the performance of each kernel by examining latency, throughput, pipeline efficiency, and memory access patterns across varying context lengths. The results, summarized in Tables V and VI, reveal significant architectural bottlenecks for quadratic-time algorithms when deployed on the NPU.
As shown in Table V, the latency of standard Causal Attention Masking scales quadratically with the sequence length ($N$). This leads to a dramatic drop in throughput, with Causal processing only 4 operations per second at a context of 8192. In contrast, sub-quadratic methods like TSA and Linear Attention maintain low latency and high throughput, demonstrating their suitability for long-context applications on this hardware.
TABLE V: Latency and throughput scaling at short ($N=512$) and long ($N=8192$) contexts.
@@ III-G Performance Scaling and Bottleneck Analysis
The underlying cause of this performance degradation is revealed in Table VI. At a context length of 8192, Causal and Retentive attention mechanisms suffer from extremely high pipeline stall rates (96.7% and 94.8%, respectively). This indicates that the NPU pipeline is mostly idle, waiting for data to be fetched from memory. The pull stage, responsible for data retrieval, is the primary contributor to these stalls. This is further corroborated by the poor cache efficiency scores for these models (7.7% for Causal and 28.1% for Retentive), which signify a low degree of data reuse and frequent, costly access to main memory. In contrast, the more structured access patterns of Linear and Toeplitz attention lead to much higher cache efficiency and significantly fewer pipeline stalls.
TABLE VI: Efficiency metrics at long context lengths. Stall and cache values are percentages; reuse is in milliseconds.
@@
III-H Impact of State Dimension
To assess the impact of model parameters on performance, we benchmarked several kernels at a fixed context length ($N=4096$) while increasing the state dimension ($d_{\text{state}}$) from the default of 16 to 128. As detailed in Table VII, increasing the state dimension leads to a predictable rise in latency across all tested models. Notably, Fourier is the most sensitive, with its latency increasing by over $10\times$, highlighting its high computational cost with respect to the state size. Linear and Toeplitz also show increased latency but remain significantly more efficient, confirming that their performance is a function of both sequence length and model dimension.
TABLE VII: Latency impact of increasing state dimension ($d_{\text{state}}$) at fixed context length $N=4096$.
@@ III-H Impact of State Dimension
III-I Cross-Class Performance Analysis
Table VIII compares representative operators from each class at N=4096, revealing fundamental performance-bottleneck relationships.
TABLE VIII: Operator class comparison at N=4096. DPU utilization directly predicts latency efficiency.
Class Operator Bottleneck DPU% SHAVE% Latency (ms)
Convolution Conv1D Direct DPU 99.6 0.0 54.71
Conv1D Dilated DPU 99.6 0.0 74.85
Conv1D FFT SHAVE 41.0 58.8 304.64
SSM SSM Sequential DPU 100.0 0.0
SSM Parallel DPU 54.8 35.9 7.75
Attention Semiseparable DPU 85.9 13.2 26.87
Toeplitz DPU 70.0 0.0* 0.59
Retentive SHAVE 28.1 71.9 39.52
Fourier DMA 48.4 0.3 45.69
Conv1D FFT measured at N=512 due to excessive latency
SSM Sequential evaluated only to N=1024
*Toeplitz fuses element-wise ops into convolution kernel
Convolution Superiority: Direct convolution operators achieve the highest DPU utilization (99.6%) by mapping exactly to the spatial MAC array’s native 2D convolution primitive. The im2col transformation converts 1D temporal convolution into spatial 2D operations that exploit full PE array parallelism. This results in predictable, near-optimal latency: 54.71 ms at N=4096 for Conv1D Direct represents just 1.12× overhead versus Toeplitz (0.59 ms) despite processing 4× longer contexts than Toeplitz’s efficient range.
SSM Efficiency-Parallelism Tradeoff: Sequential SSM provides the cleanest architectural mapping—pure DPU operations with zero SHAVE or DMA overhead—but sacrifices temporal parallelism through recurrent dependencies. Parallel SSM inverts this tradeoff: CumSum enables parallel execution but introduces 36% SHAVE occupancy at long contexts. The result is 10× faster inference (2.53 ms vs 26.38 ms at N=1024) at the cost of reduced DPU efficiency. This demonstrates a fundamental NPU design principle: parallelism gains outweigh unit utilization when bottlenecks are avoided.
Attention Variants Span Full Bottleneck Spectrum: The six attention operators cover all possible NPU bottleneck modes. Toeplitz and Semiseparable achieve DPU-dominance through structured sparsity and low-rank decomposition, respectively. Retentive becomes SHAVE-bound due to per-element decay masks. Fourier transitions to DMA-bound as FFT tensor management saturates memory bandwidth. Full Causal exhibits memory-bound behavior through cache inefficiency. This diversity suggests that attention mechanism design offers the richest optimization space for NPU co-design.
Bottleneck Predictability: A striking pattern emerges: operators with >>85% DPU utilization (Conv1D Direct/Dilated, Semiseparable) exhibit predictable, near-linear latency scaling. Operators with significant SHAVE occupancy (Retentive, Conv1D FFT) show super-linear degradation as vector cores saturate. DMA-bound operators (Fourier at N>>2048) exhibit bifurcated behavior—efficient at short contexts, catastrophic beyond memory bandwidth limits. This predictability validates our decomposition-based bottleneck analysis.
@@ IV-B Operator Intensity Characterization
TABLE IX: Operational intensity and measured performance at $N=4096$, $d_{h}=64$ (16-bit precision).
@@ IV-C Roofline Analysis
Figure 7: Roofline Modeling for Causal Inference Operators on NPU.
@@ IV-E Efficiency Analysis
TABLE X: Hardware utilization metrics at $N=4096$.
@@
This model proves that memory access patterns, not theoretical FLOP counts, dominate NPU performance. Operators must be co-designed with: (1) spatial-compatible dataflow, (2) predictable memory access, and (3) minimized DMA transfers to approach effective hardware limits.
@@ V Discussion
SHAVE Core Bottlenecks in Element-wise Operations While NPUs excel at matrix multiplication via spatial MAC arrays, element-wise operations (e.g., $\operatorname{softmax}$, $\operatorname{scaling}$) execute on general-purpose SHAVE cores. At $N>1024$, these operations dominate latency in recurrent attention variants (up to 76% utilization) with kernel fusion or acceleration being essential.
DMA Management for Memory-Intensive Ops Tensor concatenation and state management consume 40-50% of cycles in Fourier attention (Table III). DMA overheads stem from frequent allocation/deallocation of large buffers. Offloading these operations to the CPU reduces latency by 32% in tests.
@@ VI-A Long-Context Inference on Edge Platforms
Prior work has explored deploying transformer-based causal large language models (LLMs) on edge platforms [zhang2024edgeshardefficientllminference], including hardware-specific optimizations for ARM CPUs [10.1145/3700410.3702126] and FPGA-based execution through frameworks like llama.cpp [llamacpp, haris2024designing]. While these efforts target on-device inference, they are not designed for long-context scenarios and do not address the associated memory and compute bottlenecks that emerge in attention-heavy models. Our work addresses this gap by empirically analyzing a range of causal inference mechanisms—including standard transformers and structured state-space models (SSMs)—under long-context settings. This enables us to derive architectural insights that inform the co-design of attention mechanisms for Neural Processing Units (NPUs), supporting more efficient hardware-aware deployment strategies.
@@ VI-B Acceleration of Sequence Models on NPUs
Several efforts have investigated transformer acceleration on NPUs [xu2025fast, zhu2025edge], typically through operator-level scheduling or compiler-level block partitioning. However, these approaches fall short in capturing the fine-grained resource behavior required for efficient long-context inference. Other work has focused on optimizing SSMs for NPUs [das2025xamba, aalishah2025mambalitesr], leveraging architectural properties such as linear recurrence and memory compression. While successful within their respective domains, these strategies are not directly applicable to transformer-style causal attention. In contrast, our approach uses execution profiling and performance modeling—grounded in structured operator variants—to analyze architectural trade-offs across attention and SSM-style models. By leveraging structured state-space duality (SSD), we characterize how causal operators interact with NPU memory and compute hierarchies, enabling more informed co-design for future inference systems.
@@ VII Conclusion
Deploying long-context AI on edge platforms centers on resolving the fundamental mismatch between NPU architectures—optimized for dense, regular computations—and the memory-intensive, irregular access patterns of quadratic attention. Our analysis reveals catastrophic hardware under-utilization in standard approaches, while demonstrating that structured sub-quadratic operators (Toeplitz, Linear) transform the bottleneck into manageable bandwidth constraints. This necessitates a paradigm shift: throughput gains come not from incremental attention optimizations, but from co-designing causal operators that respect spatial dataflows and memory hierarchies. By aligning algorithmic structure with NPU execution models, we unlock the path to pervasive, private, and powerful edge AI.
Acknowledgment
This work was supported by the U.S. National Science Foundation (NSF) under grant CCF-1912680, the DEVCOM Army Research Lab (ARL) under grant W911NF-242-0194, and the Semiconductor Research Corporation (SRC) under grant 2024-AH-3207. We are grateful to Deepak Mathaikutty for his insightful perspectives in the development of this paper.